OpenStudy (falco276):

HELP!!! :( Given that f(x) = 2x + 5 and g(x) = x − 7, solve for f(g(x)) when x = −3.

OpenStudy (falco276):

I honestly know NOTHING ABOUT ALGEBRA 2!!!!!! I'm now taking a pretest!

OpenStudy (falco276):

@amistre64 @abb0t @ash2326 @hba @Hero @Luigi0210 @Mr.ClayLordMath @Preetha @satellite73 @thomaster

OpenStudy (falco276):

@phi

OpenStudy (luigi0210):

Plug the functions into one another

OpenStudy (phi):

the first step is find g(x) g(x) = x -7 think of it as a "rule" g(-3) means replace x with -3 in the "rule": -3 - 7 now simplify: g(-3) = -3 - 7 = ? what do you get ?

OpenStudy (cggurumanjunath):